40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13033     Entry time: Fri Jun 2 01:22:50 2017     In reply to: 13032     Reply to this: 13048
Author: gautam 
Type: Update 
Category: ASS 
Subject: ASS restoration work 

I started by checking if shaking an optic in pitch really moves it in pitch - i.e. how much PIT to YAW coupling is there. The motivation being if we aren't really dithering the optics in orthogonal DoFs, the demodulated error signals carry mixed information which the dither alignment servos get confused by. First, I checked with a low frequency dither (~4Hz) and looked at the green transmission on the video monitors. The spot seemed to respond reasonably orthogonally to both pitch and yaw excitations on either ITMY or ETMY. But looking at the Oplev control signal spectra, there seems to be a significant amount of cross coupling. ITMY YAW, ETMY PIT, and ETMY YAW have the peak in the orthogonal degree of freedom at the excitation frequency roughly 20% of the height of the DoF being driven. But for ITMY PIT, the peaks in the orthogonal DoFs are almost of equal height. This remains true even when I changed the excitation frequencies to the nominal dither alignment servo frequencies.

I then tried to see if I could get parts of the ASS working. I tried to manually align the ITM, ETM and TTs as best as I could. There are many "alignment references" - prior to the coil driver board removal, I had centered all Oplevs and also checked that both X and Y green beams had nominal transmission levels (~0.4 for GTRY, ~0.5 for GTRX). Then there are the Transmon QPDs. After trying various combinations, I was able to get good IR transmission, and reasonable GTRY.

Next, I tried running the ASS loops that use error signals demodulated at the ETM dither frequencies (so actuation is on the ITM and TT1 as per the current output matrix which I did not touch for tonight). This worked reasonably well - Attachment #1 shows that the servos were able to recover good IR transmission when various optics in the Y arm were disturbed. I used the same oscillator frequencies as in the existing burt snapshot. But the amplitudes were tweaked.

Unfortunately I had no luck enabling the servos that demodulate the ITM dithers.

The plan for daytime work tomorrow is to check the linearity of the error signals in response to static misalignment of some optics, and then optimize the elements of the output matrix.

I am uploading a .zip file with Sensoray screen-grabs of all the test-masses in their best aligned state from tonight (except ITMX face, which for some reason I can't grab).

And for good measure, the Oplev spot positions - Attachment #3.


While Gautam is working the restoration of Yarm ASS, I worked on Xarm.


Attachment 1: ASS_Y_recovery.png  31 kB  | Hide | Hide all
Attachment 2: ASS_Repairs.zip  274 kB
Attachment 3: OLs.png  8 kB  | Hide | Hide all
ELOG V3.1.3-